# A Multimodal Manufacturing Safety Chatbot: Knowledge Base Design, Benchmark Development, and Evaluation of Multiple RAG Approaches

### v1

@@ Abstract
Ensuring worker safety remains a critical challenge in modern manufacturing environments. Industry 5.0 reorients the prevailing manufacturing paradigm toward more human-centric operations. Using a design science research methodology, we identify three essential requirements for next-generation safety training systems: high accuracy, low latency, and low cost. We introduce a multimodal chatbot powered by large language models that meets these design requirements. The chatbot uses retrieval-augmented generation to ground its responses in curated regulatory and technical documentation. To evaluate our solution, we developed a domain-specific benchmark of expert-validated question and answer pairs for three representative machines: a Bridgeport manual mill, a Haas TL-1 CNC lathe, and a Universal Robots UR5e collaborative robot. We tested 24 RAG configurations using a full-factorial design and assessed them with automated evaluations of correctness, latency, and cost. Our top 2 configurations were then evaluated by ten industry experts and academic researchers. Our results show that retrieval strategy and model configuration have a significant impact on performance. The top configuration (selected for chatbot deployment) achieved an accuracy of 86.66%, an average latency of 10.04 seconds, and an average cost of \$0.005 per query. Overall, our work provides three contributions: an open-source, domain-grounded safety training chatbot; a validated benchmark for evaluating AI-assisted safety instruction; and a systematic methodology for designing and assessing AI-enabled instructional and immersive safety training systems for Industry 5.0 environments.
@@ 1 Introduction
Manufacturing continues to play a central role in the U.S. economy. In 2024, it generated more than \$2.9 trillion in annual GDP, or roughly 10% of total U.S. economic output, according to the Bureau of Economic Analysis [bea2024manufacturing]. The sector employs over 12.7 million workers [bea2025employment] and drives around 54% of the research and development conducted by American businesses with 10 or more domestic employees in 2021 [nsf2024rd]. These figures highlight both the economic importance of the sector and the need to support the safety and productivity of its workforce.
During the past 15 years, manufacturing has experienced a significant shift with the rise of smart manufacturing technologies, often referred to as Industry 4.0. These advances include the Industrial Internet of Things (IIoT), cyber-physical systems (CPS), collaborative robots (cobots), machine learning (ML), and other digital tools that provide real-time data for process monitoring and system adaptation [xu2021industry]. Even with these improvements, Industry 4.0 still faces limitations. Many systems operate close to their technological ceiling, where adding more automation produces only small gains. Legacy equipment also makes it difficult to achieve interoperability and reconfigure systems quickly. As production demands change and human robot interaction becomes more common, issues such as ergonomic strain, operator fatigue, and safety concerns continue to appear. These ongoing challenges show that productivity gains alone are not enough to ensure resilience. More recently, Industry 5.0 has emerged to address these concerns by placing human centricity, resilience, and sustainability on equal footing with efficiency [xu2021industry, lu2022outlook]. By encouraging closer collaboration between humans and intelligent machines, Industry 5.0 points toward manufacturing systems that are more adaptable, safer, and better able to recover from disruptions while continuing to improve through digital intelligence.
Despite its economic importance, manufacturing continues to face persistent safety challenges. According to the U.S. Bureau of Labor Statistics [bls2024injuries], the sector reported over 320,000 workplace injuries, making it one of the most hazardous occupational domains. Common incidents include cuts, crush injuries, and caught-in or between accidents. These events endanger workers and impose substantial economic costs on employers and the broader economy. The manufacturing sector spends an estimated \$7.5 billion each year on direct medical payments and lost wages [LibertyMutual2025WSI]. Non-fatal workplace injuries also result in an average of 11 lost workdays per employee, equivalent to approximately \$1,590 in lost productivity per case [peterson2021average].
@@
This is where the application of AI can have a positive influence. For example, prior studies have identify applications of AI in manufacturing to support operators in performing their tasks, including the presentation of maintenance guidelines, process monitoring, predicting machine downtime, and worker training [kiangala2024experimental, mayyas2020fenceless]. Modernizing worker training is a critical area for advancing worker performance and mitigating safety risk. Traditional training has often involved classroom or online lecture-based content presented upon job entry and/or at fixed intervals (e.g., annually). This approach is static and fails to reflect the complexity of modern manufacturing environments or worker experiences. In this static setup, when an operator needs information about a piece of equipment, they may not be able to easily access information in a user manual or online. There is thus a pressing need for adaptive, context-aware training systems capable of providing personalized, accurate, and readily accessible safety information in real time.
Following a design science research (DSR) methodology [peffers2007design], our work identifies three essential requirements that a next-generation manufacturing safety training solution must satisfy: high accuracy, low cost, and low latency. To address these requirements, we propose an open-source Retrieval Augmented Generation (RAG) based chatbot grounded in regulatory documents such as OSHA standards and OEM manuals. Prior work in construction safety shows that chatbot-based training can improve hazard awareness, especially for participants with limited on-site experience [zhu2022can]. One advantage identified in that work is the ability to tailor training materials to specific worker needs. By combining large language models (LLMs) with curated, domain-specific sources, our RAG system provides precise and contextually relevant safety guidance while maintaining affordability and responsiveness.
@@ 2 Literature Review
Conversational AI frameworks have begun to help operators interpret real-time industrial data. For example, Jeon et al. jeon2025chatcnc observe that limited digital literacy among shop-floor workers often makes conventional monitoring interfaces difficult to use. Their system, ChatCNC, addresses this issue by allowing operators to ask plain-language questions about machine conditions, such as whether a machine is overheating, instead of navigating complex software menus and interfaces.
@@
Research in this space reflects a broader shift toward AI-mediated instruction and monitoring that aims to reduce human error and improve safety outcomes. Natural language techniques such as LLMs are particularly attractive because they blend general reasoning capability with a conversational interface that aligns well with Industry 5.0’s emphasis on human-centric smart manufacturing. For example, Wang et al. wang2024llm developed a collaborative robot with vision–language abilities that enable it to visually map its surroundings and follow spoken or written instructions through an LLM-based parser. Similarly, Lou et al. lou2025large built an LLM-powered cognitive agent that plans and makes decisions for assembly and disassembly tasks. Their agent demonstrates context awareness and outperforms traditional automation by leveraging the LLM’s extensive knowledge and reasoning abilities. These examples show how LLMs can function as conversational intermediaries that translate operator intentions into complex technical actions.
@@
A major challenge for LLMs in manufacturing is the gap between broad model training and the specific, real-time knowledge needed on the factory floor. Retrieval-Augmented Generation (RAG) has emerged as a practical solution to this issue by pairing an LLM with a retrieval component that supplies relevant domain information such as documents, sensor readings, or knowledge base entries. Fan et al. fan2025MaViLa illustrate this approach with MaViLa, a vision–language model capable of interpreting live visuals and grounding its guidance with retrieved manufacturing knowledge. While MaViLa focuses on broad manufacturing tasks such as process understanding and skill acquisition, our work targets manufacturing safety and uses RAG to ground responses in a curated safety knowledge base.
@@
Another important development in manufacturing involves the growing adoption of multimodal interfaces that integrate vision, language, and other data modalities. Many shop-floor scenarios are inherently visual, such as identifying hazards or verifying proper use of protective equipment, which means an AI assistant that can both see and converse offers clear advantages over text-only systems. The industrial metaverse described by Li et al. li2024industrial envisions manufacturing environments where workers and AI interact through virtual and augmented reality tools that combine digital avatars, spatial audio, mixed-reality visuals, and text. These trends demonstrate that an effective manufacturing chatbot must be multimodal so that it can process images or sensor data and generate language that supports safety monitoring and operator assistance. Multimodality therefore serves as a core design principle in our work.
@@
Digital twins and virtual environments have also emerged as key platforms for integrating LLMs, RAG, and multimodal interfaces in realistic, safety-oriented settings. A digital twin is a dynamic virtual representation of a physical system that maintains continuous two-way communication with its real-world counterpart, enabling real-time monitoring, predictive simulation, control, and what-if analysis. In manufacturing, these technologies support proactive safety management and training without interrupting production. Gautam et al. gautam2025iiot demonstrate this potential by linking an LLM-based agent system with a factory digital twin. Their system uses specialized LLM agents for machine expertise, data visualization, and fault diagnosis to process streaming IIoT data and reflect the factory state in a virtual model. The digital twin then offers an interactive 3D interface where operators can query an avatar about system status or troubleshooting steps, allowing them to gain insights without approaching hazardous equipment.
@@
Taken together, these developments show the main components of a modern manufacturing safety assistant. They demonstrate that LLMs support natural communication, that RAG provides timely and relevant domain knowledge, and that multimodal understanding is essential for interpreting factory conditions. Based on these insights, we created a retrieval-augmented multimodal chatbot for manufacturing safety training and hazard prevention. Our system uses an extensive safety knowledge base, perceives the manufacturing environment, and engages in dialog to deliver practical safety guidance and real-time risk analysis.
@@ 3 Research Methodology
We address the challenge of modernizing hazard-recognition training in manufacturing by adopting a design science research methodology (DSR) [peffers2007design]. This methodological approach is well-suited for the development and evaluation of information technology solutions (artifacts) that aim to solve complex, real-world problems. Following DSR ensures that our work systematically integrates scientific rigor with practical relevance, producing outcomes that are both theoretically grounded and operationally viable within industrial settings.
In particular, our research process follows the six-stage model outlined by Peffers et al. peffers2007design, which we elaborate on throughout the paper: (1) problem identification and motivation, (2) definition of solution objectives, (3) artifact design and development, (4) demonstration, (5) evaluation, and (6) communication. We have already identified and motivated the problem in the preceding section, and the last stage of the DSR methodology, i.e., communication, is represented by this paper. That said, we now turn to the second stage, namely defining the solution objectives, which are derived “from the problem definition and knowledge of what is possible and feasible[peffers2007design]. To operationalize this stage, we established a set of design requirements that any proposed artifact must satisfy to address the underlying hazard-recognition and safety-training challenges.
@@
Our second design requirement concerns system latency, reflecting the need for timely and responsive feedback within dynamic manufacturing environments. In industrial contexts, hazards can emerge and evolve in seconds, and delayed or sluggish responses can render safety guidance ineffective or even dangerous. A training system that fails to deliver instructions, alerts, or assessments in real time risks losing its relevance to the unfolding situation on the shop floor. Therefore, any proposed solution must minimize latency in both data processing and user feedback, ensuring that workers receive immediate, context-appropriate information as they interact with machinery or simulated environments. Low latency is essential not only for maintaining user engagement and situational awareness but also for reinforcing procedural learning under realistic temporal conditions. The following definition summarizes our second design requirement.
@@
Our third design requirement concerns the cost of deployment and operation, emphasizing the importance of economic accessibility for broad adoption across the manufacturing sector. Safety innovation cannot achieve its intended societal impact if the associated technologies are prohibitively expensive to implement, maintain, or scale. Many small and medium-sized manufacturers, which account for 98 percent of U.S. manufacturing firms [sba2025facts], face constrained capital and “do not have the resources” to invest in advanced technologies and associated training [armstrong2021advanced]. Therefore, training solutions must minimize recurring expenses such as software licensing, hardware requirements, and data-processing costs, without compromising instructional quality or safety integrity.
@@
In the following sections, we design, demonstrate, and evaluate an artifact that satisfies the above design requirements. We do so by proposing the six-phase framework in Figure 1 for building manufacturing-safety chatbots grounded in expert-curated operational and safety documents. These phases align perfectly with our DSR methodology. For example, Phases 1 and 2 are part of the third DSR stage called “artifact design and development.” In particular, in Phase 1, we curate and preprocess safety materials from OSHA, the National Institute for Occupational Safety and Health (NIOSH), and Original Equipment Manufacturer (OEM) manuals to build a reliable, structured corpus. In Phase 2, we construct multiple knowledge bases that integrate keyword, semantic, and graph indices to support flexible information retrieval for question answering (Q&A). In Phase 3, we transition to the fourth design science stage, namely artifact demonstration. In particular, we demonstrate how a grounded, multimodal chatbot can be developed based on the knowledge base and indexing from the previous phase. Phases 4 to 6 are linked to the fifth stage of our DSR methodology, namely artifact evaluation. Phase 4 develops a benchmark Q&A dataset, with expert-validated questions and reference answers. We evaluate multiple retrieval-augmented generation (RAG) configurations using automated metrics in Phase 5 to identify a reasonable subset of configurations for human evaluation. In Phase 6, we employ blind human evaluation to select the “best” configuration. The evaluations in Phases 4 to 6 enable us to implement a single configuration and deploy a multimodal chatbot that integrates text, visual, and speech interactions, ensuring that its responses remain grounded in our curated source documents.
Phase 1: Source Curation and Corpus Development Phase 2: Knowledge Base and Indexing Phase 3: Multimodal Chatbot Phase 4: Benchmark Q&A Dataset Development Phase 5: Automated Evaluation and Experimental Design Phase 6: Human Evaluation [  Collect safety documents @  Extract and clean text Content segmentation V  Filename standardization Cloud-Based keyword and semantic search indices `  Local index for keyword search Graph-based indices Ψ  Retrieval routing across heterogeneous indices Chat UI Multimodal LLM for text and vision capabilities ?  Speech capabilities Θ  Unified multimodal user interface v Draft safety questions from collected sources Write gold answers Validate with industry and safety experts 8  Finalize reference answers Θ  Define experimental factors and their levels Generate answers for each configuration Compute automated metrics Z  Rank and select top $n$ configurations O  Curate human evaluation Q&A benchmark set 6  Conduct blind human grading of curated set y  Aggregate human scores per configuration Select the final and winning configuration

### v2

@@ Abstract
Ensuring worker safety remains a critical challenge in modern manufacturing environments. Industry 5.0 reorients the prevailing manufacturing paradigm toward more human-centric operations. Using a design science research methodology, we identify three essential requirements for next-generation safety training systems: high accuracy, low latency, and low cost. We introduce a multimodal chatbot powered by large language models that meets these design requirements. The chatbot uses retrieval-augmented generation to ground its responses in curated regulatory and technical documentation. To evaluate our solution, we developed a domain-specific benchmark of expert-validated question and answer pairs for three representative machines: a Bridgeport manual mill, a Haas TL-1 CNC lathe, and a Universal Robots UR5e collaborative robot. We tested 24 RAG configurations using a full-factorial design and assessed them with automated evaluations of correctness, latency, and cost. Our top 2 configurations were then evaluated by ten industry experts and academic researchers. Our results show that retrieval strategy and model configuration have a significant impact on performance. The top configuration, selected for chatbot deployment, achieved an accuracy of 86.66%, an average cost of \$0.005 per query, and an average end-to-end latency of 10.04 seconds. This latency is practical for delivering a complete safety instruction and is measured from query submission to full instruction delivery rather than generation onset. Overall, our work provides three contributions: an open-source, domain-grounded safety training chatbot; a validated benchmark for evaluating AI-assisted safety instruction; and a systematic methodology for designing and assessing AI-enabled instructional and immersive safety training systems for Industry 5.0 environments.
@@ 1 Introduction
Manufacturing continues to play a central role in the U.S. economy. In 2024, it generated more than \$2.9 trillion in annual GDP, or roughly 10% of total U.S. economic output, according to the Bureau of Economic Analysis [Bureau of Economic Analysis, 2025]. The sector employs over 12.7 million workers [U.S. Bureau of Economic Analysis, 2025] and drives around 54% of the research and development conducted by American businesses with 10 or more domestic employees in 2021 [National Center for Science and Engineering Statistics, 2024]. These figures highlight both the economic importance of the sector and the need to support the safety and productivity of its workforce.
During the past 15 years, manufacturing has experienced a significant shift with the rise of smart manufacturing technologies, often referred to as Industry 4.0. These advances include the Industrial Internet of Things (IIoT), cyber-physical systems (CPS), collaborative robots (cobots), machine learning (ML), and other digital tools that provide real-time data for process monitoring and system adaptation [Xu et al., 2021]. Even with these improvements, Industry 4.0 still faces limitations. Many systems operate close to their technological ceiling, where adding more automation produces only small gains. Legacy equipment also makes it difficult to achieve interoperability and reconfigure systems quickly. As production demands change and human robot interaction becomes more common, issues such as ergonomic strain, operator fatigue, and safety concerns continue to appear. These ongoing challenges show that productivity gains alone are not enough to ensure resilience. More recently, Industry 5.0 has emerged to address these concerns by placing human centricity, resilience, and sustainability on equal footing with efficiency [Xu et al., 2021, Lu et al., 2022]. By encouraging closer collaboration between humans and intelligent machines, Industry 5.0 points toward manufacturing systems that are more adaptable, safer, and better able to recover from disruptions while continuing to improve through digital intelligence.
[…]
@@
This is where the application of AI can have a positive influence. For example, prior studies have identify applications of AI in manufacturing to support operators in performing their tasks, including the presentation of maintenance guidelines, process monitoring, predicting machine downtime, and worker training [Kiangala and Wang, 2024, Mayyas et al., 2020]. Modernizing worker training is a critical area for advancing worker performance and mitigating safety risk. Traditional training has often involved classroom or online lecture-based content presented upon job entry and/or at fixed intervals (e.g., annually). This approach is static and fails to reflect the complexity of modern manufacturing environments or worker experiences. In this static setup, when an operator needs information about a piece of equipment, they may not be able to easily access information in a user manual or online. There is thus a pressing need for adaptive, context-aware training systems capable of providing personalized, accurate, and readily accessible safety information in real time.
Following a design science research (DSR) methodology [Peffers et al., 2007], our work identifies three essential requirements that a next-generation manufacturing safety training solution must satisfy: high accuracy, low cost, and low latency. To address these requirements, we propose an open-source, Retrieval Augmented Generation (RAG) based chatbot grounded in regulatory documents such as OSHA standards and OEM manuals. Prior work in construction safety shows that chatbot-based training can improve hazard awareness, especially for participants with limited on-site experience [Zhu et al., 2022]. One advantage identified in that work is the ability to tailor training materials to specific worker needs. By combining large language models (LLMs) with curated, domain-specific sources, our RAG system provides precise and contextually relevant safety guidance while maintaining affordability and responsiveness.
@@ 2 Literature Review
Conversational AI frameworks have begun to help operators interpret real-time industrial data. For example, Jeon et al. Jeon et al. [2025] observe that limited digital literacy among shop-floor workers often makes conventional monitoring interfaces difficult to use. Their system, ChatCNC, addresses this issue by allowing operators to ask plain-language questions about machine conditions, such as whether a machine is overheating, instead of navigating complex software menus and interfaces.
@@
Research in this space reflects a broader shift toward AI-mediated instruction and monitoring that aims to reduce human error and improve safety outcomes. Natural language techniques such as LLMs are particularly attractive because they blend general reasoning capability with a conversational interface that aligns well with Industry 5.0’s emphasis on human-centric smart manufacturing. For example, Wang et al. Wang et al. [2024] developed a collaborative robot with vision–language abilities that enable it to visually map its surroundings and follow spoken or written instructions through an LLM-based parser. Similarly, Lou et al. Lou et al. [2025] built an LLM-powered cognitive agent that plans and makes decisions for assembly and disassembly tasks. Their agent demonstrates context awareness and outperforms traditional automation by leveraging the LLM’s extensive knowledge and reasoning abilities. These examples show how LLMs can function as conversational intermediaries that translate operator intentions into complex technical actions.
@@
A major challenge for LLMs in manufacturing is the gap between broad model training and the specific, real-time knowledge needed on the factory floor. RAG has emerged as a practical solution to this issue by pairing an LLM with a retrieval component that supplies relevant domain information such as documents, sensor readings, or knowledge base entries. Fan et al. Fan et al. [2025] illustrate this approach with MaViLa, a vision–language model capable of interpreting live visuals and grounding its guidance with retrieved manufacturing knowledge. While MaViLa focuses on broad manufacturing tasks such as process understanding and skill acquisition, our work targets manufacturing safety and uses RAG to ground responses in a curated safety knowledge base.
@@
Another important development in manufacturing involves the growing adoption of multimodal interfaces that integrate vision, language, and other data modalities. Many shop-floor scenarios are inherently visual, such as identifying hazards or verifying proper use of protective equipment, which means an AI assistant that can both see and converse offers clear advantages over text-only systems. The industrial metaverse described by Li et al. Li et al. [2024] envisions manufacturing environments where workers and AI interact through virtual and augmented reality tools that combine digital avatars, spatial audio, mixed-reality visuals, and text. These trends demonstrate that an effective manufacturing chatbot must be multimodal so that it can process images or sensor data and generate language that supports safety monitoring and operator assistance. Multimodality therefore serves as a core design principle in our work.
@@
Digital twins and virtual environments have also emerged as key platforms for integrating LLMs, RAG, and multimodal interfaces in realistic, safety-oriented settings. A digital twin is a dynamic virtual representation of a physical system that maintains continuous two-way communication with its real-world counterpart, enabling real-time monitoring, predictive simulation, control, and what-if analysis. In manufacturing, these technologies support proactive safety management and training without interrupting production. Gautam et al. Gautam et al. [2025] demonstrate this potential by linking an LLM-based agent system with a factory digital twin. Their system uses specialized LLM agents for machine expertise, data visualization, and fault diagnosis to process streaming IIoT data and reflect the factory state in a virtual model. The digital twin then offers an interactive 3D interface where operators can query an avatar about system status or troubleshooting steps, allowing them to gain insights without approaching hazardous equipment.
@@
Taken together, these developments show the main components of a modern manufacturing safety assistant. They demonstrate that LLMs support natural communication, that RAG provides timely and relevant domain knowledge, and that multimodal understanding is essential for interpreting factory conditions. Despite this progress, the literature still lacks (1) domain-grounded, safety-focused benchmarks with expert-validated gold answers, (2) systematic, controlled comparisons of multiple retrieval strategies and model configurations using accuracy–latency–cost criteria, and (3) a reproducible methodology linking knowledge-base design choices to end-to-end RAG performance for manufacturing safety guidance. Our work addresses these gaps by proposing a RAG-based chatbot for manufacturing safety training and hazard prevention. Our system uses an extensive safety knowledge base, perceives the manufacturing environment, and engages in dialog to deliver practical safety guidance and real-time risk analysis. Moreover, we contribute an expert-validated benchmark spanning representative manufacturing equipment and describe a systematic methodology for selecting RAG configurations.
@@ 3 Research Methodology
We address the challenge of modernizing hazard-recognition training in manufacturing by adopting a design science research methodology (DSR) [Peffers et al., 2007]. This methodological approach is well-suited for the development and evaluation of information technology solutions (artifacts) that aim to solve complex, real-world problems. Following DSR ensures that our work systematically integrates scientific rigor with practical relevance, producing outcomes that are both theoretically grounded and operationally viable within industrial settings.
In particular, our research process follows the six-stage model outlined by Peffers et al. Peffers et al. [2007], which we elaborate on throughout the paper: (1) problem identification and motivation, (2) definition of solution objectives, (3) artifact design and development, (4) demonstration, (5) evaluation, and (6) communication. We have already identified and motivated the problem in the preceding section, and the last stage of the DSR methodology, i.e., communication, is represented by this paper. That said, we now turn to the second stage, namely defining the solution objectives, which are derived “from the problem definition and knowledge of what is possible and feasible[Peffers et al., 2007]. To operationalize this stage, we established a set of design requirements that any proposed artifact must satisfy to address the underlying hazard-recognition and safety-training challenges.
@@
Our second design requirement concerns system latency, reflecting the need for timely and responsive feedback within dynamic manufacturing environments. In industrial contexts, hazards can emerge and evolve in seconds, and delayed or sluggish responses can render safety guidance ineffective or even dangerous. A training system that fails to deliver instructions, alerts, or assessments in real time risks losing its relevance to the unfolding situation on the shop floor. Therefore, any proposed solution must minimize end-to-end latency in both data processing and user feedback, ensuring that workers receive immediate, context-appropriate information as they interact with machinery or simulated environments. Here, latency is defined as the total time from user query submission to complete instruction delivery (i.e., full response rendered), not time-to-first-token or generation onset. Low end-to-end latency is essential not only for maintaining user engagement and situational awareness but also for reinforcing procedural learning under realistic temporal conditions. The following definition summarizes our second design requirement.
@@
Our third design requirement concerns the cost of deployment and operation, emphasizing the importance of economic accessibility for broad adoption across the manufacturing sector. Safety innovation cannot achieve its intended societal impact if the associated technologies are prohibitively expensive to implement, maintain, or scale. Many small and medium-sized manufacturers, which account for 98 percent of U.S. manufacturing firms [Office of Advocacy, 2025], face constrained capital and “do not have the resources” to invest in advanced technologies and associated training [Armstrong et al., 2021]. Therefore, training solutions must minimize recurring expenses such as software licensing, hardware requirements, and data-processing costs, without compromising instructional quality or safety integrity.
@@
In the following sections, we design, demonstrate, and evaluate an artifact that satisfies the above design requirements. We do so by proposing the six-phase framework in Figure 1 for building manufacturing-safety chatbots grounded in expert-curated operational and safety documents. These phases align perfectly with our DSR methodology. For example, Phases 1 and 2 are part of the third DSR stage called “artifact design and development.” In particular, in Phase 1, we curate and preprocess safety materials from OSHA, the National Institute for Occupational Safety and Health (NIOSH), and Original Equipment Manufacturer (OEM) manuals to build a reliable, structured corpus. In Phase 2, we construct multiple knowledge bases that integrate keyword, semantic, and graph indices to support flexible information retrieval for question answering (Q&A). In Phase 3, we transition to the fourth design science stage, namely artifact demonstration. In particular, we demonstrate how a grounded, multimodal chatbot can be developed based on the knowledge base and indexing from the previous phase. Phases 4 to 6 are linked to the fifth stage of our DSR methodology, namely artifact evaluation. Phase 4 develops a benchmark Q&A dataset, with expert-validated questions and reference answers. We evaluate multiple RAG configurations using automated metrics in Phase 5 to identify a reasonable subset of configurations for human evaluation. In Phase 6, we employ blind human evaluation to select the “best” configuration. The evaluations in Phases 4 to 6 enable us to implement a single configuration and deploy a multimodal chatbot that integrates text, visual, and speech interactions.
Phase 1: Source Curation and Corpus Development Phase 2: Knowledge Base and Indexing Phase 3: Multimodal Chatbot Phase 4: Benchmark Q&A Dataset Development Phase 5: Automated Evaluation and Experimental Design Phase 6: Human Evaluation Collect safety documents Extract and clean text Content segmentation Filename standardization Cloud-Based keyword and semantic search indices Local index for keyword search Graph-based indices Retrieval routing across heterogeneous indices Chat UI Multimodal LLM for text and vision capabilities Speech capabilities Unified multimodal user interface Draft safety questions from collected sources Write gold answers Validate with industry and safety experts Finalize reference answers Define experimental factors and their levels Generate answers for each configuration Compute automated metrics Rank and select top $n$ configurations Curate human evaluation Q&A benchmark set Conduct blind human grading of curated set Aggregate human scores per configuration Select the final and winning configuration